OpenStudy (anonymous):

Suppose that f satisfies f(x+y)=f(x)+f(y), and that f is continuous at 0. Prove that f is continuous on R.

OpenStudy (anonymous):

what is the definition of continuous for all \(a\in \mathbb R\) you have to show that \[\lim_{x\to a}f(x)=f(a)\] now \[\lim_{x\to a}f(x)=\lim{h\to 0}f(a+h)=f(a)+\lim_{h\to 0}f(h)=f(a)+f(0)\] the first equality by a change of variable, the second by the condition that \(f(x+y)=f(x)+f(y)\) and the third by the fact that \(f\) is continuous at \(x=0\) all that is left for you show is that if \(f(x+y)=f(x)+f(y)\) then it must be true that \(f(0)=0\)

OpenStudy (anonymous):

i will leave that up to you, but if you want a hint, start with \[f(0)=f(0+0)\]

OpenStudy (anonymous):

okay, i will try first, but if i don't know to start with, i will ask you again ya..